If the perimeter of a rectangle is 36 ft and the width is 10 ft, find its length. [Hint: The perimeter of a rectangle is given by P = 2L + 2W.]
Answer by checkley71(8403)

36=2L+2*10
2L=36-20
2L=16
L=16/2
L=8
PROOF
36=2*8+2+10
36=16+20
36=36